# Self-organizing railway traffic management

### v1

@@ VI Experimental setup
To assess the performance of the proposed SO-TMS and compare it against baseline approaches, we set up a closed loop experimental framework that connects the OpenTrack microscopic railway simulator [27] with the TMS through an API that regulates information exchange between between the two programs, as well as synchronization. The OpenTrack simulator and the TMS exchange information exploiting the RTTP and TS format are explained in [7] and exemplified in [6]. We consider a control area corresponding to the $60$ km line connecting Segrate and Ospitaletto, in Italy, as represented in Figure 3. Two lines branch off from the Pioltello station: the Pioltello–Treviglio–Rovato (VE LL) line which from Pioltello reaches Rovato, encountering a large number of stations and stops; the Pioltello–Bivio Casirate–Treviglio (VE AVAC) line which runs parallel to the VE LL line up to Melzo Scalo, and then enters the Milan-Brescia highspeed line at the Casirate junction. From Casirate, through an interconnection, the line also branches off to Treviglio. About $150$ trains per day circulate in this control area. Each train can use between $2$ and 64 51264\,512 alternative routes, with a mean and median of 9 3669\,366 and $256$, respectively.
@@
Here, freight, intercity, regional and high speed trains belonging to different companies share the same tracks. No detailed information is available on the historic train loads, be them passenger or freight trains, nor on the company operating each train. To mimic the competitive context in place, we assume that each train is independent from the others. Each train knows its own cost of delay, and makes an educated guess on the one of the other trains. This cost is the weight used in the objective function for the hypothesis generation process (Section V-B). In agreement with the industrial partners of the SORTEDMOBILITY project [35], weights depend on the types of trains: for freight trains, weights vary in the interval $[15,24]$; for regional trains, in the interval $[21,36]$; for intercity trains, in the interval $[19,27]$; for high-speed trains, in the interval $[24,40]$; for empty rides of passenger trains, in the interval $[15,18]$.
@@
Here, each colored line is a train path, the horizontal lines correspond to stations along the line. Dashed lines show the path of trains using the VA AVAC branch in Figure 3, between Bivio Casirate and Treviglio. To make instances more tractable, we partition each scheduled train in the timetable into a sequence of shorter trains: the shorter trains use the same rolling stock, and they are linked by re-utilization constraints at stations. In particular, a single train going from origin to destination through multiple stops is represented by multiple trains doing just one stop: one going from the origin to the first stop, one from the first to the second stop, and so on, up to the last one going from the last stop to the destination. Between pairs of subsequent trains in these sequences, compatibility constraints state that the TDS where the first train arrives must be the same from where the second train departs, and that the second train cannot depart before the first one has arrived plus a minimum rolling stock reutilization time, which here corresponds to the minimum dwell time. The so-obtained timetable is characterized by a significantly higher number of trains, but with a reduced number of alternative routes: it includes $456$ trains having between $1$ and $44$ alternative routes, with mean and median of $6$ and $4$, respectively.
@@
In our experiments, we consider $35$ perturbed traffic scenarios that include 93 trains. These are all trains entering between 1:00 and 4:00 AM in the compressed timetable. Looking at the original timetable, this time span covers all trains entering the control area before 3:15 PM. In each perturbed scenario, trains enter the infrastructure with a random delay. The distributions from which entrance delays are drawn are deduced from historical data. They are piecewise uniform distributions, summarized in Table I. They depend on the type of train and on the entrance point. In the table, we indicate the probability that each train has a delay in the interval heading each column. Depending on the perturbation, the total duration of scenarios exploiting SO-TMS varies between $4$ hours and $30$ minutes and $6$ hours: the exit time of the last train from the control area determines the end of the scenario, and this depends on the specific trains that suffer delays. In the closed loop, decision making by any TMS is triggered every five minutes. Depending on the duration of each scenario, with SO-TMS decisions are made between $56$ and $74$ times. When triggering decision making, a TS is produced, including all focal trains and all trains or entering the control area within the next $40$ minutes.
With this framework, we compare the performance of three traffic management approaches:
@@
  • CEN: the decision making is carried out centrally through the RECIFE-MILP algorithm [29];
@@
In SO, given the current TS, the neighborhood selection uses a time horizon of $T_{h}=15$ minutes, and each train generates at most two hypotheses: the optimal one, or the best one the hypothesis generation found within the time limit, and the one corresponding to the current RTTP. The time limit for each run of the hypothesis generation is set to three minutes. Hypothesis generation considers the weighted sum of the delays as the objective function, and the generated hypotheses are shared among neighboring trains. Finally, the consensus process runs for a maximum number of 100 000100\ 000 iterations, after which a consensus failure would be reported. In CEN, the optimization is carried out minimizing the sum of delays, hence the weights are not considered as we assume this is private information available only to trains. For each iteration of the optimization process, the same time limit of three minutes is considered to solve instances including all trains in the TS. This CEN setup mimics the one of previous studies on this type of deployment [4, 31]. In FCFS, a myopic strategy is enacted, so that whenever a conflict arises, the train that can pass first has priority, and the others follow.
@@ VII Experimental results
In this section, we analyze the performance of the proposed TMS, comparing the results obtained with SO, CEN and FCFS. The metric we use to assess the performance of SO is the percentage improvement with respect to either baseline approach, i.e., CEN or FCFS. In the following, we will refer to it as improvement, for brevity. We compute this improvement for both total weighted and unweighted delay over the $93$ trains of each scenario. Specifically, we consider the delay with which they exit the control area in simulation, with respect to the timetable. In the computation of the weighted delay, we use the weights that are considered by trains in their own hypothesis generation in SO. Hence, this total weighted delay represents the actual total cost of delay.
@@
Figure 5 reports the distribution of the improvement of SO over both CEN and FCFS across all scenarios. It is clear that SO outperforms both CEN and FCFS: all points of the distributions are strictly above zero, meaning that SO never produces higher delays than the other approaches.
@@
Figure 5: Percentage improvement of total weighted and unweighted delay of SO w.r.t CEN (left) and FCFS (right).
The average improvements are represented by diamonds. In the SO vs. CEN comparison, the average improvement is about $5\%$ for the weighted delay, and about $4\%$ for the unweighted one. For SO vs FCFS, these values rise to $31\%$ and $28\%$, respectively. All improvements are statistically significant according to the Wilcoxon signed-rank test with $\alpha=0.05$.
@@
The very large improvements observed with respect to FCFS show that the scenarios are actually quite complex, and optimization allows a much more effective traffic management than the application of the myopic rule. However, the optimization-based approaches do not achieve the same traffic effectiveness, SO being always slightly better than CEN and reaching an improvement of around $20\%$ in the best case, regardless of the type of delay considered.
@@
There are multiple factors that could explain how this improvement can be achieved. First, we analyze the impact of the different objective functions used for SO and CEN, that is, the unweighted versus weighted delay. Indeed, CEN considers the former, while SO considers the latter in each train’s hypothesis generation. Although the merge repair operation described in Section V-E is performed centrally in SO, and as such considers unweighted delays in the objective function, we consider the impact of this operation minimal, also considering that all repair actions are re-assessed by trains at later stages, thus considering weighted delays. For this analysis, we study the conflict resolutions decided by both SO and CEN at each iteration and for each scenario. The explicit information of what conflicts are dealt with when making decisions is not available, as the solution generation procedure employed is based on branch and bound and does not specifically identify conflicts. Hence, we define a proxy for the set of conflict resolutions decided, which we refer to as quasi-conflicts. At each iteration, given the RTTP produced by the considered TMS (either CEN or SO), we identify all pairs of trains in which one or both exit the control area with a positive delay, and which have paths that immediately follow each other on at least one TDS. If the paths of a pair of trains follow each other immediately on more than a TDS, we consider the first one. These situations correspond to the quasi-conflicts we use as a proxy for the set of conflicts solved by the TMS. In the $35$ scenarios, CEN returns RTTPs with $517$ quasi-conflicts in average (between $359$ and $747$), while SO with $424$ (between $336$ and $526$). The pairs of trains involved in these quasi-conflicts are at most $140$ for CEN and $132$ for SO, with averages across scenarios of $121$ and $110$. The higher quantities observed for CEN are consistent with the worse performance observed for this approach in terms of total delay, be it weighted or unweighted: although there is no strictly monotonic relation, often more conflicts (and therefore quasi-conflicts) result in higher delays. When we look more closely at these quasi-conflicts, we see that, in average, both CEN and SO make the train with the higher weight pass first in $46\%$ of these cases. The train with the lower weight passes first in $48\%$ of the cases for CEN and in $49\%$ of the cases for SO, in average, the remaining quasi-conflicts being between trains with equal weight. From these numbers, it emerges that SO does not appear to favor higher-weight trains. Indeed, in these scenarios, the use of different weights for train delay in the objective function does not seem to impact SO choices to the point of having it inverting train passing orders in many conflicts. Overall, we can conclude that the improvement of SO over CEN does not depend on the different objective functions used.
@@
Another relevant factor that differentiates SO and CEN is problem decomposition. Indeed, in the former, trains generate hypotheses that possibly modify only the paths of their neighbors. They hence operate on sub-instances of the rtRTMP, while CEN considers all paths as modifiable and operates on complete instances. On the one hand, this intuitively makes it easier for SO to find the optimal solution for each traffic management sub-instance trains happen to encounter. But, on the other hand, the choices made could be myopic, as they do not consider possible conflicts in the long run. In particular, in SO, trains may make decisions that are good regarding the first trains they may interact with, but that may result not to be good later on.
@@
Figure 6: Representation of traffic and its management evolution across scenarios and iterations: thick lines represent average values across scenarios while thin lines represent individual scenarios. A: Total number of trains. B: Number of focal trains. C: Average size of the neighborhood of a train. D: Percentage optimality gap per iteration, for RECIFE-MILP in CEN and its hypothesis generation variant in SO. E: Number of connected components. F: percentage of focal trains included in the largest connected component.
@@
Figure 6 shows how the traffic and its management evolve over time. During the simulated period, the number of trains considered in the traffic management (i.e., those that are in the control area or will enter it in a time window of 40 minutes) reaches a maximum value of about $40$ trains (see Figure 6A). Of these trains, in average, at most $\sim 25$ are focal: they are present in the control area or will enter it within the neighborhood selection time horizon ($T_{h}=15$ minutes) and, as such, they generate hypotheses and actively participate to the traffic management process in SO (see Figure 6B). For each focal train, the neighborhood size is rather small, with $\sim 10$ trains in average, suggesting that problem decomposition can actually reduce the size of the instances (see Figure 6C). After iteration 35 (around 4:00 AM), these numbers start to decrease as this is the end of the considered horizon, no more trains enter the control area and the ones that are present progressively exit it. Iterations between 0 and 35 are referred to as the regime interval, while iterations beyond this interval are referred to as the tail of the simulated scenario. The fact that train neighborhoods are rather small, even in the regime interval, leads to a higher quality of the produced solutions. Figure 6D reports the percentage optimality gap achieved by RECIFE-MILP in CEN and by its hypothesis generation variant in SO along iterations: this is the difference between the value of the best solution found and the lower bound computed during the branch and bound procedure. Clearly, SO benefits from problem decomposition, achieving optimality in almost all cases, while CEN is often penalized by the difficulty of the overall traffic situation and typically returns sub-optimal solutions with gaps of almost $20\%$ in average.
@@
The impact of myopia due to problem decomposition is very difficult to quantify, as the fact of making decisions periodically in closed-loop partially mitigates it. As a proxy for the impact of myopic decisions at each iteration, though, we look at the infeasibilities that arise after the merge operations, which require an intervention to repair the RTTP. Our analysis reveals that such repair procedures is required in only $3.5\%$ of the cases. This indicates that SO effectively manages traffic even in the long run, thanks to the consideration of the paths of trains out of the neighborhoods as hard constraints. This reduces the need for centralized intervention only to a handful of cases. Another element possibly impacting the myopia of decisions is the number and size of the connected components of $G_{I}$ (see Figure 6E and  6F). Indeed, if all trains were part of the same connected component of the graph, their decisions would be all consistent, at least to some extent. In our experiments, we observe that in the regime interval there often is only one large connected component, involving almost all the trains present in the control area or entering it within $T_{h}$. When more than one connected component exists and the maximum size of these components is around $100\%$, it means that one of the components is actually made of one or two trains, while the other includes all the remaining focal trains. The number of connected components increases up to five only in the tail, when trains start to exit the control area. We then consider, for all perturbation scenarios, the relationship between the number of connected components and the occurrence of repair procedures, limited to the iterations in which the repair procedure was necessary. In other words, given all the iterations in which at least one perturbation scenario resulted in a final repair procedure, we compute the number of connected components for all perturbation scenarios, distinguishing between those where repair was necessary and those where it was not. Here, a Mann-Whitney U test does not support a statistically significant difference between the number of connected components in instances requiring merge repair and those that do not ($p\text{-value}=0.77$). This suggests that the graph connectivity is not the primary driver for merge repair interventions, which instead could depend on the degree of nodes in the interaction graph combined with the designed hypothesis compatibility, that limits the detection of conflicts only for pairs involving the focal trains. Measures to address this issue will be further investigated in future work.
@@
In summary, problem decomposition appears to be a definite strength for SO with respect to CEN, as no remarkable effect of myopic decision making is observed.
@@
To conclude the analysis of the results, we observe the performance of the consensus process designed for hypothesis selection. First of all, trains reached consensus on all the $C_{T}=1974$ calls of the hypothesis selection module ($100\%$). Figure 7A shows the distribution of regret, i.e., the distribution of the difference between the total delay of the selected hypotheses with respect to the optimal one. It emerges that in $C_{O}=1783$ cases ($90.3\%$), the consensus converges to the optimal solution, while, in the remaining cases, the maximum regret is only $0.6\%$. This signifies that even when the consensus solution is not optimal, it is very close to the optimal one. The average number of decisions per train required to reach consensus follows the distribution shown in Figure 7B. In $C_{I}=1614$ cases ($81.8\%$), consensus is reached immediately, requiring no iterative adjustments. Given that trains initiate the hypothesis selection process by selecting their best hypothesis, an immediate consensus signifies that no hypothesis change was necessary: all selected hypotheses are compatible. Hence, the problem decomposition directly leads to a feasible solution. When further decision steps are necessary, the trains converge to a solution in at most $8$ steps, in average. Out of the $C_{T}-C_{I}=360$ cases in which at least one iteration of the hypothesis selection was required, $C_{O}-C_{I}=169$ cases achieved the optimal solution (i.e., null regret). Hence, nearly half of the times ($\frac{C_{O}-C_{I}}{C_{T}-C_{I}}=47\%$), executing the consensus process achieved optimality.

### v2

@@ VI Experimental setup
To assess the performance of the proposed SO-TMS and compare it against baseline approaches, we set up a closed loop experimental framework that connects the OpenTrack microscopic railway simulator [35] with the TMS through an API that regulates the information exchanges between the two programs, as well as synchronization. The OpenTrack simulator and the TMS exchange information exploiting the RTTP and TS formats as explained in [13] and illustrated in [12]. Specifically, the TS lists the trains that either are currently in the control area (in-area trains) or are predicted to enter it (out-of-area trains) within a fixed horizon from the current time (40 minutes in this paper). For each in-area train, the TS indicates which TDS the train’s head is currently occupying and the timestamp at which it was first detected there. For each out-of-area train, the TS indicates the expected entry timestamp. The RTTP reports, for each train, the sequence of TDSs that the train is scheduled to traverse and the time at which it is predicted to do so. From these times, passing orders can be obtained. In our implementation, the data encoded in TS and RTTP are exchanged via xml files.
We consider a control area corresponding to the $60$ km line connecting Segrate and Ospitaletto Travagliato, in Italy, as represented in Figure 3. Two lines branch off from the Pioltello station: the Pioltello–Treviglio–Rovato (VE LL) line which from Pioltello reaches Rovato, encountering a large number of stations and stops; the Pioltello–Bivio Casirate–Treviglio (VE AVAC) line which runs parallel to the VE LL line up to Melzo Scalo, and then enters the Milan-Brescia highspeed line at the Casirate junction. From Casirate, through an interconnection, the line also branches off to Treviglio. About $150$ trains per day circulate in this control area. Each train can use between $2$ and 64 51264\,512 alternative routes, with a mean and median of 9 3669\,366 and $256$, respectively.
@@
Here, freight, intercity, regional and high speed trains belonging to different companies share the same tracks. No detailed information is available on the historic train loads, be them passenger or freight trains, nor on the company operating each train. To mimic the competitive context, we assume that each train is independent from the others. Each train knows its own cost of delay, and makes an educated guess on the one of the other trains. This cost is the weight used in the objective function for the hypothesis generation process (Section V-B). The choice of having all independent trains does not precisely capture reality, where many trains are operated by the same companies. However, we conjecture that this does not have a major impact on the results, as we assume trains aim at operating in an overall well-performing system by considering all trains’ delays in their objective functions. In agreement with the industrial partners of the SORTEDMOBILITY project [45], weights depend on the types of trains: for freight trains, weights vary in the interval $[15,24]$; for regional trains, in the interval $[21,36]$; for intercity trains, in the interval $[19,27]$; for high-speed trains, in the interval $[24,40]$; for empty rides of passenger trains, in the interval $[15,18]$. From these intervals, we randomly select the weight for each train. We consider a uniform probability distribution inspired by what has been done in the literature for assigning capacity to competing railway companies during timetable production [4].
@@
Here, each colored line is a train path, the horizontal lines correspond to stations along the line. Dashed lines show the path of trains using the VA AVAC branch in Figure 3, between Bivio Casirate and Treviglio. To make instances more tractable, we partition each scheduled train in the timetable into a sequence of shorter trains: the shorter trains use the same rolling stock, and they are linked by re-utilization constraints at stations. In particular, a single train going from origin to destination through multiple stops is represented by multiple trains doing just one stop: one going from the origin to the first stop, one from the first to the second stop, and so on, up to the last one going from the last stop to the destination. Between pairs of subsequent trains in these sequences, compatibility constraints state that the TDS where the first train arrives must be the same from where the second train departs, and that the second train cannot depart before the first one has arrived plus a minimum rolling stock reutilization time, which here corresponds to the minimum dwell time. The so-obtained timetable is characterized by a significantly higher number of trains, but with a reduced number of alternative routes: it includes $456$ trains having between $1$ and $44$ alternative routes, with mean and median of $6$ and $4$, respectively. Partitioning the original trains makes the instance more tractable for RECIFE-MILP, which defines one binary variable for each train-route pair. Indeed, when trains pass stations with multiple alternative platforms, their number of routes grows combinatorially. Instead, partitioned trains that pass through fewer stations have fewer routes, and the number of variables grows at a much slower pace. For example, a train passing four stations, each with three platforms, will require 81 binary variables, while only 27 after being partitioned (three trains with nine routes each). In the neighborhood selection, partitioned trains coming from the same original one are considered jointly: as soon as one of them is in a neighborhood, all the others are also included. By doing so, the only impact of the partitioning is the speed-up of RECIFE-MILP.
@@
In our experiments, we consider $35$ perturbed traffic scenarios that include 93 trains. These are all trains entering between 1:00 and 4:00 AM in the compressed timetable. Looking at the original timetable, this time span covers all trains entering the control area before 3:15 PM. In each perturbed scenario, trains enter the infrastructure with a random delay, and any additional change to the plan is a consequence of the entrance delays and the resulting traffic congestion (e.g., the need to reduce train speeds or to change dwell times at stations). Such entrance delay perturbations are very frequently considered in the literature on the rtRTMP [38, 7, 9]. The distributions from which entrance delays are drawn are deduced from historical data. They are piecewise uniform distributions, summarized in Table I. They depend on the type of train and on the entrance point. In the table, we indicate the probability that each train has a delay in the interval heading each column. For each train, we draw a random number between zero and one, and we identify the piecewise distribution corresponding to its type and entrance point. If the number drawn is smaller than or equal to the end of the first step (e.g., 0.24 for a passenger train from Segrate), then we assign no entrance delay to the train. If it is between the beginning and the end of the second step, we assign a delay equal to the corresponding value in this step. For example, if the random number drawn for a passenger train from Segrate is 0.475, we first identify the corresponding step, i.e., the second one that corresponds to a delay between zero and five minutes. Then we observe that it falls right in the middle of this step (0.4750.24=0.2370.475-0.24=0.237, which is $0.47/2$), and we assign a delay equal to the middle of the corresponding interval to the train (2.5 minutes).
Depending on the perturbation, the total duration of scenarios exploiting SO-TMS varies between $4$ hours and $30$ minutes and $6$ hours: the exit time of the last train from the control area determines the end of the scenario, and this depends on the specific trains that suffer delays. In the closed loop, decision making by any TMS is triggered every five minutes. Depending on the duration of each scenario, with SO-TMS decisions are made between $56$ and $74$ times. When triggering decision making, a TS is produced, including all focal trains and all trains present or entering the control area within the next $40$ minutes.
With this framework, we compare the performance of five traffic management approaches:
@@
  • CEN-1: the decision making is carried out centrally through the RECIFE-MILP algorithm [38] in a real-time context;
  • CEN-2: the decision making is carried out by using the ACO-TRSP algorithm [37] to limit the number of alternative routes available for each train, and by chosing routing and passing orders through the RECIFE-MILP algorithm [38] in a real-time context;
  • CEN*: the decision making is carried out centrally through the RECIFE-MILP algorithm [38] allowing a very long computational time, aiming to find optimal solutions;
@@
In SO, given the current TS, the neighborhood selection uses a time horizon of $T_{h}=15$ minutes, and each train generates at most two hypotheses: the optimal one, or the best one the hypothesis generation found within the time limit, and the one corresponding to the current RTTP. The time limit for each run of the hypothesis generation is set to three minutes. Hypothesis generation considers the weighted sum of the delays as the objective function, and the generated hypotheses are shared among neighboring trains. Finally, in our simulation of the decentralized hypothesis selection process, we simulate asynchronous train decisions by running sequential iterations, in each of which we randomly select one train among the focal ones to make a decision. The time limit is represented by a maximum number of iterations, which we set to 100 000100\ 000. If no feasible solution is found after these iterations, a consensus failure is reported.
In CEN-1, the optimization is carried out minimizing the sum of delays, hence the weights are not considered as we assume this is private information available only to trains. For each execution of the optimization, the same time limit of three minutes is considered to solve instances including all trains in the TS. This setup mimics the one of previous studies on this type of deployment [10, 40].
In CEN-2, we use the same setting as in CEN-1, but, at each execution, 30 seconds are used by ACO-TRSP to find the six most suitable alternative routes per train, and two minutes and 30 seconds are available for RECIEF-MILP, as in the reference paper [37].
In CEN*, we use the same setting as in CEN-1, but, at each execution, we let RECIFE-MILP run for one hour.
In FCFS, a myopic strategy is enacted, so that whenever a conflict arises, the train that can pass first has priority, and the others follow.
@@ VII Experimental results
In this section, we analyze the performance of the proposed TMS, comparing the results obtained with SO, CEN-1, CEN-2, CEN* and FCFS. Each SO module requires the following minimum, average, and maximum computational times per execution (in seconds): neighborhood selection - 5, 12 and 23; hypothesis generation - 5, 28 and 180; hypothesis compatibility - 0, 1 and 5; hypothesis selection - 1, 2 and 2; merge - 5, 6 and 32.
@@
The metric we use to assess the performance of SO is the percentage improvement with respect to each baseline approach, i.e., CEN-1, CEN-2, CEN* or FCFS. In the following, we will refer to it as improvement, for brevity. We compute this improvement for both total weighted and unweighted delay over the $93$ trains of each scenario. Specifically, we consider the delay with which they exit the control area in simulation, with respect to the timetable. In the computation of the weighted delay, we use the weights that are considered by trains in their own hypothesis generation in SO. Hence, this total weighted delay represents the actual total cost of delay.
Note that, while the total unweighted delay corresponds to CEN-1’s, CEN-2’s and CEN*’s objective function, the total weighted one is never used to optimize: in SO no train knows the exact weight of the others and can only use educated guesses. Nonetheless, we think these two indicators are appropriate for our analysis, as they are the ones that best represent the overall performance of traffic management.
@@
Figure 5 reports the distribution of the improvement of SO over the four benchmarks across all scenarios. It is clear that SO outperforms FCFS, CEN-1 and CEN-2: for the first two, all points of the distributions are strictly above zero, meaning that SO never produces higher delays than the other approaches. For CEN-2, SO is worse than this benchmark in two out of the 35 scenarios, by $0.1$ and $0.2\%$, and $0.2$ and $0.3\%$ for unweighted and weighted delay, respectively. In the remaining 33 scenarios, SO regularly improves the solution by a few percentage points.
@@
Figure 5: Percentage improvement of total weighted and unweighted delay of SO w.r.t FCFS, CEN-1, CEN-2 and CEN*.
The average improvements are represented by diamonds in the figure. In the SO vs. FCFS comparison, the average improvement is about $31\%$ for the weighted delay, and about $28\%$ for the unweighted one. Given the already good performance of the centralized benchmarks, for SO vs CEN-1, these values decrease to $5\%$ and $4\%$, respectively. Since, in these experiments, CEN-2 performs better than CEN-1 due to the large number of route alternatives in the original instances, the improvement brought by SO when compared to the former further decreases to $3\%$ and $2\%$ for the weighted and unweighted delay, in average. All improvements are statistically significant according to the Wilcoxon signed-rank test with $\alpha=0.05$. The comparison between SO and CEN* shows that the former, in the short computational time it is allowed to use, typically gets extremely close to the optimal solution, or to the best solution found by the state-of-the-art RECIFE-MILP in a one-hour execution. Indeed, this time is not always enough to find the optimal solution, as some traffic situations result in extremely difficult instances: in $81\%$ out of the total 2150 executions, CEN* fails to prove solution optimality, and gaps remain between $0.1\%$ and $13.2\%$ ($0.5\%$ in average). In all scenarios, there is at least one execution in which optimality is not proven. In $11$ scenarios, SO manages to deal with the difficult instances more effectively, improving over CEN* in both weighted and unweighted delay. According to the Wilcoxon signed-rank test with $\alpha=0.05$, the difference remains statistically significant in favor of CEN* when unweighted delay is considered. Instead, no significant difference is observed in terms of weighted delay.
@@
The very large improvements observed with respect to FCFS show that the scenarios are actually quite complex, and optimization allows a much more effective traffic management than the application of the myopic rule. However, all optimization-based approaches do not achieve the same traffic effectiveness, SO being extremely close to CEN* and almost always slightly better than both CEN-1 and CEN-2 and reaching an improvement of around $20\%$ in the best case, regardless of the type of delay considered.
@@
There are multiple factors that could explain how this improvement can be achieved. First, we analyze the impact of the different objective functions used for SO and CENs, that is, the unweighted versus weighted delay. Indeed, CENs consider the former, while SO considers the latter in each train’s hypothesis generation, even if the weights of the trains different from the focal one are not precisely known. Although the merge repair operation described in Section V-E is performed centrally in SO, and as such minimizes unweighted delays in the objective function, we consider the impact of this operation minimal, also knowing that all repair actions are re-assessed by trains at later stages, thus based on weighted delays. For this analysis, we study the conflict resolutions decided by both SO and CEN-1 at each execution and for each scenario. The explicit information of what conflicts are dealt with when making decisions is not available, as the solution generation procedure employed is based on branch and bound and does not specifically identify conflicts. Hence, we define a proxy for the set of conflict resolutions decided, which we refer to as quasi-conflicts. At each execution, given the RTTP produced by the considered TMS (either CEN-1 or SO), we identify all pairs of trains in which one or both exit the control area with a positive delay, and which have paths that immediately follow each other on at least one TDS. If the paths of a pair of trains follow each other immediately on more than a TDS, we consider the first one. These situations correspond to the quasi-conflicts we use as a proxy for the set of conflicts solved by the TMS. In the $35$ scenarios, CEN-1 returns RTTPs with $517$ quasi-conflicts in average (between $359$ and $747$), while SO with $424$ (between $336$ and $526$). The pairs of trains involved in these quasi-conflicts are at most $140$ for CEN-1 and $132$ for SO, with averages across scenarios of $121$ and $110$. The higher quantities observed for CEN-1 are consistent with the worse performance observed for this approach in terms of total delay, be it weighted or unweighted: although there is no strictly monotonic relation, often more conflicts (and therefore quasi-conflicts) result in higher delays. When we look more closely at these quasi-conflicts, we see that, in average, both CEN-1 and SO make the train with the higher weight pass first in $46\%$ of these cases. The train with the lower weight passes first in $48\%$ of the cases for CEN-1 and in $49\%$ of the cases for SO, in average, the remaining quasi-conflicts being between trains with equal weight. From these numbers, it emerges that SO does not appear to favor higher-weight trains. Indeed, in these scenarios, the use of different weights for train delay in the objective function does not seem to impact SO choices to the point of having it inverting train passing orders in many conflicts. Overall, we can conclude that the improvement of SO over CEN-1 does not depend on the different objective functions used.
@@
Another relevant factor that differentiates SO and CENs is problem decomposition. Indeed, in the former, trains generate hypotheses that possibly modify only the paths of their neighbors. They hence operate on sub-instances of the rtRTMP, while CENs considers all paths as modifiable and operates on complete instances. On the one hand, this intuitively makes it easier for SO to find the optimal solution for each traffic management sub-instance trains happen to encounter. But, on the other hand, the choices made could be myopic, as they do not consider possible conflicts in the long run. In particular, in SO, trains may make decisions that are good regarding the first trains they may interact with, but that may result not to be good later on.
@@
Figure 6: Representation of traffic and its management evolution across scenarios and executions: thick lines represent average values across scenarios while thin lines represent individual scenarios. A: Total number of trains. B: Number of focal trains. C: Average size of the neighborhood of a train. D: Percentage optimality gap per execution, for RECIFE-MILP in CEN-1 and 2, and CEN* and its hypothesis generation variant in SO. E: Number of connected components. F: percentage of focal trains included in the largest connected component.
@@
Figure 6 shows how the traffic and its management evolve over time. During the simulated period, the number of trains considered in the traffic management (i.e., those that are in the control area or will enter it in a time window of 40 minutes) reaches a maximum value of about $40$ trains (see Figure 6A). Of these trains, in average, at most $\sim 25$ are focal: they are present in the control area or will enter it within the neighborhood selection time horizon ($T_{h}=15$ minutes) and, as such, they generate hypotheses and actively participate to the traffic management process in SO (see Figure 6B). For each focal train, the neighborhood size is rather small, with $\sim 10$ trains in average, suggesting that problem decomposition can actually reduce the size of the instances (see Figure 6C). After execution 35 (around 4:00 AM), these numbers start to decrease as this is the end of the considered horizon, no more trains enter the control area and the ones that are present progressively exit it. Executions between 0 and 35 are referred to as the regime interval, while executions beyond this interval are referred to as the tail of the simulated scenario. The fact that train neighborhoods are rather small, even in the regime interval, leads to a higher quality of the produced solutions. Figure 6D reports the percentage optimality gap achieved by RECIFE-MILP in CEN-1, 2 and CEN*, and by its hypothesis generation variant in SO along executions: this is the difference between the value of the best solution found and the lower bound computed during the branch and bound procedure. Clearly, SO benefits from problem decomposition, achieving optimality in almost all cases, while CENs are often penalized by the difficulty of the overall traffic situation and typically returns sub-optimal solutions with gaps of almost $20\%$ in average for CEN-1, and $4\%$ for CEN-2 that operates on a smaller search space thanks to the initial reduction of the number of alternative routes through ACO-TRSP. CEN*, despite its much longer computation time, registers an average gap of $0.5\%$.
The impact of myopia due to problem decomposition is very difficult to quantify, as the fact of making decisions periodically in closed-loop partially mitigates it. As a proxy for the impact of myopic decisions at each execution, though, we look at the infeasibilities that arise after the merge operations, which require an intervention to repair the RTTP. Our analysis reveals that such repair procedures is required in only $3.5\%$ of the cases. This indicates that SO effectively manages traffic even in the long run, thanks to the consideration of the paths of trains out of the neighborhoods as hard constraints in the hypothesis generation module. This reduces the need for centralized intervention only to a handful of cases, in which trains not identified as potentially in conflict in the neighborhood selection module change their path in ways that create inconsistencies in the far-away future (see Section V-E).
@@
Another element possibly impacting the myopia of decisions is the number and size of the connected components of $G_{I}$ (see Figure 6E and  6F). Indeed, if all trains were part of the same connected component of the graph, their decisions would be all consistent, at least to some extent. In our experiments, we observe that in the regime interval there often is only one large connected component, involving almost all the trains present in the control area or entering it within $T_{h}$. When more than one connected component exists and the maximum size of these components is around $100\%$, it means that one of the components is actually made of one or two trains, while the other includes all the remaining focal trains. The number of connected components increases up to five only in the tail, when trains start to exit the control area. We then consider, for all perturbation scenarios, the relationship between the number of connected components and the occurrence of repair procedures, limited to the executions in which the repair procedure was necessary. In other words, given all the executions in which at least one perturbation scenario resulted in a final repair procedure, we compute the number of connected components for all perturbation scenarios, distinguishing between those where repair was necessary and those where it was not. Here, a Mann-Whitney U test does not support a statistically significant difference between the number of connected components in instances requiring merge repair and those that do not ($p\text{-value}=0.77$). This suggests that the graph connectivity is not the primary driver for merge repair interventions, which instead could depend on the degree of nodes in the interaction graph combined with the designed hypothesis compatibility, that limits the detection of conflicts only for pairs involving the focal trains. How different graph characteristics are linked to merge repair needs will be further investigated in future work.
@@
In summary, problem decomposition appears to be a definite strength for SO with respect to CEN’s, as no remarkable effect of myopic decision making is observed.
@@
To conclude the analysis of the results, we observe the performance of the consensus process designed for hypothesis selection. First of all, trains reach consensus on all the $C_{T}=1974$ calls of the hypothesis selection module ($100\%$). Figure 7A shows the distribution of regret, i.e., the distribution of the difference between the total delay of the selected hypotheses with respect to the optimal one. It emerges that in $C_{O}=1783$ cases ($90.3\%$), the consensus converges to the optimal solution, while, in the remaining cases, the maximum regret is only $0.6\%$. This signifies that even when the consensus solution is not optimal, it is very close to the optimal one. The average number of decisions per train required to reach consensus follows the distribution shown in Figure 7B. In $C_{I}=1614$ cases ($81.8\%$), consensus is reached immediately, requiring no iterative adjustments. Given that trains initiate the hypothesis selection process by selecting their best hypothesis, an immediate consensus signifies that no hypothesis change was necessary: all selected hypotheses are compatible. Hence, the problem decomposition directly leads to a feasible solution. When further decision steps are necessary, the trains converge to a solution in at most $8$ steps, in average. Out of the $C_{T}-C_{I}=360$ cases in which at least one iteration of the hypothesis selection is required, $C_{O}-C_{I}=169$ cases achieve the optimal solution (i.e., null regret). Hence, nearly half of the times ($\frac{C_{O}-C_{I}}{C_{T}-C_{I}}=47\%$), executing the consensus process achieves optimality.
@@
For an analysis of the performance of the consensus process in abstract complex networks, see [14].